OpenStudy (anonymous):

earl opened a credit card at a department store with an APR of 20.95% compounded daily. What is the APY on this credit card? A. 5.74% B. 23.30% C. 41.90% D. 46.29%

jimthompson5910 (jim_thompson5910):

Use the formula APY = (1 + r/n)^(n) - 1 where r = APR (in decimal form) n = compounding frequency